Suppose a t-shirt shop decides to decrease the price of their most popular shirt by 10%. Their sales increase from 200 shirts per day to 220 shirts per day as a result. 


Calculate PED and determine if demand for their shirt is elastic, inelastic or unit elastic.




1
Expert's answer
2021-11-16T11:35:20-0500

"PED= \frac{Percentage Change in Quantity}{Percentage Change in Price}"

The Percentage Change in Quantity:

"=\frac{220-200}{200}\times100=10" %

Percentage Change in Price:=10%

"PED=\frac{10}{10}=1."

The demand is unit elastic. This is because any change in price leads to an equally proportion change in quantity demanded.